The average age of three men is 24 years.
A. What is the sum of their ages?
B. If two of the men are 22 years old, how old is the third?

Answer by checkley79(3341)

3*24=72 ANS.
(22+22+X)/3=24
(44+X)/3=24
44+X=24*3
44+X=72
X=72-44
X=28 AGE OF THE THIRD MAN.